# Ramification filtration via deformations, II

### v3

@@ Abstract.
Let $\mathcal{K}$ be a field of formal Laurent series with coefficients in a finite field of characteristic $p$. For $M\in\mathbb{N}$, let $\mathcal{G}_{<p,M}$ be the maximal quotient of the Galois group of $\mathcal{K}$ of period $p^{M}$ and nilpotent class $<p$ and $\{\mathcal{G}_{<p,M}^{(v)}\}_{v\geqslant 0}$ — the filtration by ramification subgroups in upper numbering. Let $\mathcal{G}_{<p,M}=G(\mathcal{L})$ be the identification of nilpotent Artin-Schreier theory: here $G(\mathcal{L})$ is the group obtained from a suitable profinite Lie $\mathbb{Z}/p^{M}$-algebra $\mathcal{L}$ via the Campbell-Hausdorff composition law. We develop new techniques to obtain a “geometrical” construction of the ideals $\mathcal{L}^{(v)}$ such that $G(\mathcal{L}^{(v)})=\mathcal{G}_{<p,M}^{(v)}$. Given $v_{0}\geqslant 1$, we construct a decreasing central filtration $\mathcal{L}(w)$, $1\leqslant w\leqslant p$, on $\mathcal{L}$, an epimorphism of Lie $\mathbb{Z}/p^{M}$-algebras $\bar{\mathcal{V}}:\bar{\mathcal{L}}^{{\dagger}}\to\bar{\mathcal{L}}:=\mathcal{L}/\mathcal{L}(p)$, and a unipotent action $\Omega$ of $\mathbb{Z}/p^{M}$ on $\bar{\mathcal{L}}^{{\dagger}}$, which induces the identity action on $\bar{\mathcal{L}}$. Suppose $d\Omega=B^{{\dagger}}$, where $B^{{\dagger}}\in\operatorname{Diff}\bar{\mathcal{L}}^{{\dagger}}$, and ¯$\bar{\mathcal{L}}^{{\dagger}[v_{0}]}$ is the ideal of $\bar{\mathcal{L}}^{{\dagger}}$ generated by the elements of $B^{{\dagger}}(\bar{\mathcal{L}}^{{\dagger}})$. Our main result states that the ramification ideal $\mathcal{L}^{(v_{0})}$ appears as the preimage of the ideal in $\bar{\mathcal{L}}$ generated by $\bar{\mathcal{V}}B^{{\dagger}}(\bar{\mathcal{L}}^{{\dagger}[v_{0}]})$. In the last section we apply this to the explicit construction of generators of ¯$\bar{\mathcal{L}}^{(v_{0})}$. The paper justifies a geometrical origin of ramification subgroups of $\Gamma_{K}$ and can be used for further developing of non-abelian local class field theory.
@@ 0.1. Motivation
For a cdvf $K$, consider the decreasing filtration by normal ramification subgroups $\{\Gamma_{K}^{(v)}\}_{v\geqslant 0}$ in $\Gamma_{K}$, [28]. This (ramification) filtration allows us to recover arithmetic invariants (say, discriminants, differentes, etc) of subextensions $L$ of $K$ in $K_{sep}$. The importance of ramification filtration was never doubted because this is quite non-trivial number theoretic interpretation of the fundamental geometric concept of ramification. However, during very long time the structure of ramification subgroups $\Gamma_{K}^{(v)}$ was known just modulo the subgroup of second commutators of $\Gamma_{K}$. (In other words, we knew only the images of $\Gamma_{K}^{(v)}$ in the maximal abelian quotient $\Gamma_{K}^{ab}$ of $\Gamma_{K}$ given by local class field theory.) Only in [26] Mochizuki proved that in the mixed characteristic case $\Gamma_{K}$ together with additional structure given by its ramification filtration behaves like a Galois group of algebraic number field. In other words, if $K$ is a cdvf of characteristic 0 then the functor $K\mapsto(\Gamma_{K},\{\Gamma_{K}^{(v)}\}_{v\geqslant 0})$ is fully faithful. This gave the first formal confirmation of importance of ramification theory in the number theoretic setting. As a result, the ramification filtration should not anymore treated just as a technical concept, it is strict invariant of $K$ and, as such, requires careful and detailed study. Notice that Mochizuki’s proof is a spectacular application of the theory of $p$-adic representations; at the same time his proof is highly non-efficient.
It is well-known that the group structure of $\Gamma_{K}$ can be easily recovered from its maximal pro-$p$-quotient $\Gamma_{K}(p)=\mathrm{Gal}(K(p)/K)$, where $K(p)$ is the maximal $p$-extension of $K$ in $K_{sep}$, cf. [18]. This group is provided with induced ramification filtration $\{\Gamma_{K}(p)^{(v)}\}_{v\geqslant 0}$. In [1, 2, 3] the author obtained first explicit description (at substantially non-abelian level) of the subgroups $\Gamma_{K}(p)^{(v)}$ modulo the subgroup $C_{p}(\Gamma_{K}(p))$ of commutators of order $\geqslant p$ in $\Gamma_{K}(p)$. This approach systematically used the equivalence of the categories of $p$-groups and Lie $\mathbb{Z}_{p}$-algebras with nilpotent class $<p$ and new techniques which allowed us to work with ramification properties in terms of the appropriate Lie algebras. These results already found the following applications.
@@ 0.1.1. Mixed characteristic case
In [9, 10, 11] we used the Fontaine-Wintenberger field-of-norms functor [33] to extend the approach from [1, 2, 3] to the case of cdvf’s $K$ of mixed characteristic under the assumption that $K$ contains non-trivial $p$-th roots of unity. In particular, suppose that $K$ contains a primitive $p^{M}$-th root of unity, $M\geqslant 1$, and let $K(s,M)$ be the maximal $p$-extension of $K$ in $K(p)$ with the Galois group of period $p^{M}$ and nilpotent class $s<p$. Then we obtained explicit value for the number $v(s,M)$ such that $\mathrm{Gal}(K(s,M)/K)^{(v)}=e$ iff $v>v(s,M)$.
@@ 0.1.2. Local analogue of the Grothendieck Conjecture
In [4, 7] we obtained quite efficient generalisation of Mochizuki’s result [26] in the context of $p$-extensions for cdvf $K$ of any characteristic. Namely, if $K$ is cdvf then the functor $K\mapsto(\Gamma_{K}(p),\{\Gamma_{K}(p)^{(v)}\}_{v\geqslant 0})$ is fully faithful. If $p\geqslant 3$ the proof of the equicharacteristic case in [4] uses explicit description of the subgroups $\Gamma_{K}(p)^{(v)}$ modulo third commutators $C_{3}(\Gamma_{K}(p))\supset C_{p}(\Gamma_{K}(p))$. Then the mixed chracteristic case is deduced via application of the field-of-norms functor. In the case $p=2$ our approach does not work because modulo $C_{p}(\Gamma_{K}(p))=C_{2}(\Gamma_{K}(2))$ we are just in the abelian situation where the ramification filtration does not carry essential information. However, in [7] we studied the images of ramification filtration in the Magnus algebra associated with $\Gamma_{K}(2)$ modulo subgroup $C_{3}(\Gamma_{K}(2))$ of third commutators. This new technique allowed us to prove the case $p=2$ in the both (equicharacteristic and mixed characteristic) cases. Our method even worked in the case of higher local fields, cf. [5, 6].
@@
Question. Describe the subgroups $\pi(\Gamma_{K}^{(v)})\subset GL_{\mathbb{Z}_{p}}(V)$.
@@ 0.1.4. Non-abelian explicit reciprocity laws (ERL)
Since $K_{\pi}/K$ is abelian all $\mathrm{Gal}(K_{\pi}/K)^{(v)}$ come from $\Gamma_{K}^{ab(v)}$ which are well-known by local class field theory. Then the corresponding images of $\Gamma_{K}^{(v)}$ can be found via one-or-another ERL for the Hilbert symbol. For example, suppose $K=k((t_{0}))$ contains a primitive $p^{M}$-th root of unity and $[K_{\pi}:K]=p^{M}$ then $K_{\pi}=K(\root p^{M} \of{v_{\pi}})$, $v_{\pi}\in K^{*}$. So the images $\pi(\Gamma_{K}^{(v)})$ can be calculated via calculations with the $M$-th Hilbert symbols $(u,v_{\pi})$, where $u\equiv 1\,\mathrm{mod}\,t_{0}^{v}$. The study of different ERL’s for the Hilbert symbol (and its analogs) is one of most popular areas of Algebraic Number Theory. It has very rich history [25] started with Kummer’s work in 1859. Known ERL’s have more-or-less similar structure: the answer appears as a residue of suitable explicitly constructed differential form. As example, cf. fundamental contributions of Shafarevich [30], Artin-Hasse [14], Iwasawa [20], Brückner-Vostokov [15, 32], Witt [17], Coates-Wiles [16], et al. If $V$ from 0.1.3 is not cyclic one-or-another non-abelian version of the Hilbert symbol and the corresponding ERL’s could be used to study the question from Section 0.1.3 in full generality. Such symbol should use as essential ingredients nilpotent versions of local class field theory and of Kummer or Artin-Schreier theory. The idea of creating nilpotent class field theory was very popular for a very long time, e.g. [30, 22, 21]. Recently, F. Laubie [23] obtained such theory by generalizing the approach from [22]. It is based on the idea of iteration of Lubin-Tate theories via the field-of-norms functor. This theory posseses many nice properties but is not adjusted for obtaining (even very well-known) explicit information about involved Galois groups. A nilpotent analog of the Artin-Schreier theory was constructed by the author in [1, 2]; it allowed us to obtain new results about ramification filtration in [1, 3]. Introduction of nilpotent analog of Hilbert symbol could be stated now as the question about linking the Laubie and our theories. This will allow us to benefit from functorial properties of the first and explicit constructions of the second theories. Notice that the case of “small” represntations [13] gives a positive feeling about existence of ERL’s in the non-abelian context.
@@ 0.1.5. Geometrical approach to ramification filtration
The above applications not only illustrate a fundamental role of explicit results from [1, 2, 3] but also rise the question about their geometric origin and more structural approach. The proofs in [1] are reasonably technical but their modulo $p^{M}$-version in [2, 3] relies on heavy computations with enveloping algebras. When working with Lie algebras of nilpotent class $<p$ the most efficient and clear approach should use one-or-another deformation theory. Such theory was recently constructed in the case of Galois groups of period $p$ in [12] and this gave us geometrical approach to [1]. In this paper we construct more general deformation theory which adjusted to the modulo $p^{M}$ situation. This theory not just gives the results of [2, 3] but also considerably simplifies the approach from [12]. We expect that this new techniques could be applied to develop the modulo $p^{M}$-case of results obtained in the mixed characteristic case in [9, 10]. Note that in [12] we gave also an explicit description of arbitrarily long segments of ramification filtration (not just of a single ramification subgroup) in the case of Galois groups of period $p$. This description admits a straight generalisation to the modulo $p^{M}$-situation. As a result, we have a substantial control on explicit information about arithmetical structure of Galois groups of cdvf’s. We also expect the existence of other fundamental relations between Arithmetic and Geometry of cdvf’s.
0.2. The content
Everywhere in the paper $M\in\mathbb{N}$ and $v_{0}\in\mathbb{R}_{>0}$ are fixed. For $s\in\mathbb{N}$ and a topological group $G$, we denote by $C_{s}(G)$ the closure of the subgroup of $G$ generated by the commutators of order $\geqslant s$. If $L$ is a toplogical Lie algebra then $C_{s}(L)$ is the closure of the ideal generated by commutators of degree $\geqslant s$. For any topological $A$-module $M$ and $B$, we use the notation $M_{B}:=M\hat{\otimes}_{A}B$.
@@ 0.2. The content
Consider a decreasing filtration by ramification subgroups in the upper numbering $\{\mathcal{G}^{(v)}_{<p,M}\}_{v\geqslant 0}$ of $\mathcal{G}_{<p,M}$. In [12] we constructed a deformation theory which allowed us to retrieve (in essentially geometrical way) the results from [1]. In this paper, we obtain a generalisation (and also a considerable simplification) of our approach from [12] to the case of Galois groups $\mathcal{G}_{<p,M}$ with arbitrary exponent $p^{M}$. We also explain how this theory can be applied to recover the structure of the ramification subgroups $\mathcal{G}_{<p,M}^{(v)}$. Our approach could be overviewed as follows.
a) Let $\mathcal{K}=k((t_{0}))$, where $k\simeq\mathbb{F}_{p^{N_{0}}}$. Then we can use the uniformiser $t_{0}$ to define the lifts modulo $p^{M}$, $O_{M}(\mathcal{K})=W_{M}(k)((t))\subset W_{M}(\mathcal{K})$ and $O_{M}(\mathcal{K}_{sep})\subset W_{M}(\mathcal{K}_{sep})$. Here $W_{M}$ is the functor of Witt vectors of length $M$ and $t=[t_{0}]\in W_{M}(\mathcal{K})$ is the Teichmuller representative of $t_{0}$. This allows us to apply the nilpotent version of Artin-Shreier theory, cf. [2], to fix an identification $\eta^{0}:\mathcal{G}_{<p,M}\simeq G(\mathcal{L})$, where $\mathcal{L}$ is explicitly constructed pro-finite Lie $\mathbb{Z}/p^{M}$-algebra. The identification $\eta^{0}$ appears via choosing a suitable $e\in\mathcal{L}_{O_{M}(\mathcal{K})}$, taking $f\in\mathcal{L}_{O_{M}(\mathcal{K}_{sep})}$ such that $\sigma(f)=e\circ f$ and setting for any $\tau\in\mathcal{G}_{<p,M}$, $\eta^{0}(\tau)=(-f)\circ\tau(f)\in G(\mathcal{L})$. This provides us also with “ramification” ideals $\mathcal{L}^{(v)}$ in $\mathcal{L}$ such that $\eta^{0}(\mathcal{G}_{<p,M}^{(v)})=\mathcal{L}^{(v)}$.
b) The extension of scalars $\mathcal{L}_{W_{M}(k)}$ appears with the set of explicit generators $\{D_{an}\ |\ \mathrm{gcd}(a,p)=1,n\in\mathbb{Z}/N_{0}\}\cup\{D_{0}\}$. We use them to define a decreasing filtration of ideals $\mathcal{L}=\mathcal{L}(1)\supset\mathcal{L}(2)\supset\ldots\supset\mathcal{L}(p)$ and provide ¯$\bar{\mathcal{L}}=\mathcal{L}/\mathcal{L}(p)$ with induced filtration ¯$\bar{\mathcal{L}}(w)$, $1\leqslant w\leqslant p$. The algebra $\bar{\mathcal{L}}$ is considerably smaller than $\mathcal{L}$; it is generated over $W_{M}(k)$ by all $D_{an}$ with $0\leqslant a<(p-1)v_{0}$. However, we prove that $\mathcal{L}(p)\subset\mathcal{L}^{(v_{0})}$ and, as a result, it will be sufficient to describe the image ¯$\bar{\mathcal{L}}^{(v_{0})}$ of $\mathcal{L}^{(v_{0})}$ in $\bar{\mathcal{L}}$.

### v5

@@ Abstract.
Let $\mathcal{K}$ be a field of formal Laurent series with coefficients in a finite field of characteristic $p$. For $M\in\mathbb{N}$, let $\mathcal{G}_{<p,M}$ be the maximal quotient of the Galois group of $\mathcal{K}$ of period $p^{M}$ and nilpotent class $<p$ and let $\{\mathcal{G}_{<p,M}^{(v)}\}_{v\geqslant 0}$ be the filtration by ramification subgroups in upper numbering. We use the identification $\mathcal{G}_{<p,M}=G(\mathcal{L})$ of nilpotent Artin-Schreier theory: here $G(\mathcal{L})$ is the group obtained from a suitable profinite Lie $\mathbb{Z}/p^{M}$-algebra $\mathcal{L}$ via the Campbell-Hausdorff composition law. We develop new techniques to obtain a “geometrical” construction of the ideals $\mathcal{L}^{(v)}$ such that $G(\mathcal{L}^{(v)})=\mathcal{G}_{<p,M}^{(v)}$. Given $v_{0}\geqslant 1$, we construct a decreasing central filtration $\mathcal{L}(w)$, $1\leqslant w\leqslant p$, on $\mathcal{L}$, an epimorphism of Lie $\mathbb{Z}/p^{M}$-algebras $\bar{\mathcal{V}}:\bar{\mathcal{L}}^{{\dagger}}\longrightarrow\bar{\mathcal{L}}:=\mathcal{L}/\mathcal{L}(p)$, and a unipotent action $\Omega$ of $\mathbb{Z}/p^{M}$ on $\bar{\mathcal{L}}^{{\dagger}}$, which induces the identity action on $\bar{\mathcal{L}}$. Suppose $d\Omega=B^{{\dagger}}$, where $B^{{\dagger}}\in\operatorname{Diff}\bar{\mathcal{L}}^{{\dagger}}$, and ¯$\bar{\mathcal{L}}^{{\dagger}[v_{0}]}$ is the ideal of $\bar{\mathcal{L}}^{{\dagger}}$ generated by the elements of $B^{{\dagger}}(\bar{\mathcal{L}}^{{\dagger}})$. Our main result states that the ramification ideal $\mathcal{L}^{(v_{0})}$ appears as the preimage of the ideal in $\bar{\mathcal{L}}$ generated by $\bar{\mathcal{V}}B^{{\dagger}}(\bar{\mathcal{L}}^{{\dagger}[v_{0}]})$. In the last section we apply this to the explicit construction of generators of ¯$\bar{\mathcal{L}}^{(v_{0})}$. The paper justifies a geometrical origin of ramification subgroups of $\Gamma_{\mathcal{K}}$ and can be used for further developing of non-abelian local class field theory.
@@ 0.1. Motivation
For a cdvf $K$, consider the decreasing filtration by normal ramification subgroups $\{\Gamma_{K}^{(v)}\}_{v\geqslant 0}$ in $\Gamma_{K}$, [29]. This (ramification) filtration allows us to recover arithmetic invariants (say, discriminants, differentes, etc) of subextensions $L$ of $K$ in $K_{sep}$. The importance of ramification filtration has been never doubted because this is highly non-trivial number theoretic interpretation of the fundamental geometric concept of ramification. However, during very long time the structure of ramification subgroups $\Gamma_{K}^{(v)}$ was known just modulo the subgroup of second commutators of $\Gamma_{K}$. (In other words, we knew only the images of $\Gamma_{K}^{(v)}$ in the maximal abelian quotient $\Gamma_{K}^{ab}$ of $\Gamma_{K}$ given by local class field theory.) Only in [27] Mochizuki proved that in the mixed characteristic case the group $\Gamma_{K}$ together with additional structure given by its ramification filtration behaves like a Galois group of algebraic number field. In other words, if $K$ is a cdvf of characteristic 0 then the functor $K\mapsto(\Gamma_{K},\{\Gamma_{K}^{(v)}\}_{v\geqslant 0})$ is fully faithful. This gave the first formal confirmation of the importance of ramification theory in the number theoretic setting. As a result, the ramification filtration should be not anymore treated just as a technical concept, it is a strict invariant of $K$ and, as such, requires very careful and detailed study. Notice that Mochizuki’s proof is a spectacular application of the theory of $p$-adic representations; at the same time his proof is highly non-efficient.
It is well-known that the group structure of $\Gamma_{K}$ can be easily recovered from its maximal pro-$p$-quotient $\Gamma_{K}(p)=\mathrm{Gal}(K(p)/K)$, where $K(p)$ is the maximal $p$-extension of $K$ in $K_{sep}$, cf. [19]. This group is provided with induced ramification filtration $\{\Gamma_{K}(p)^{(v)}\}_{v\geqslant 0}$. In [1, 2, 3] the author obtained first explicit description (at substantially non-abelian level) of the subgroups $\Gamma_{K}(p)^{(v)}$ modulo the subgroup $C_{p}(\Gamma_{K}(p))$ of commutators of order $\geqslant p$ in $\Gamma_{K}(p)$. This approach systematically used the equivalence of the categories of $p$-groups and Lie $\mathbb{Z}_{p}$-algebras with nilpotent class $<p$. It is based on new techniques which allowed us to work with ramification properties in terms of the appropriate Lie algebras. These results already found the following applications.
@@ 0.1.1. Mixed characteristic case
In [9, 10, 11] we used the Fontaine-Wintenberger field-of-norms functor [34] to extend the approach from [1, 2, 3] to the case of cdvf’s $K$ of mixed characteristic under the assumption that $K$ contains non-trivial $p$-th roots of unity. In particular, suppose that $K$ contains a primitive $p^{M}$-th root of unity, $M\geqslant 1$, and let $K(s,M)$ be the maximal $p$-extension of $K$ in $K(p)$ with the Galois group of period $p^{M}$ and nilpotent class $s<p$. Then we obtained explicit value for the number $v(s,M)$ (the last jump) such that $\mathrm{Gal}(K(s,M)/K)^{(v)}=e$ iff $v>v(s,M)$.
@@ 0.1.2. Local analogue of the Grothendieck Conjecture
In [4, 7] we obtained by essentially different method quite efficient generalisation of Mochizuki’s result [27] in the context of $p$-extensions for cdvf $K$ of any characteristic. Namely, if $K$ is a cdvf then the functor $K\mapsto(\Gamma_{K}(p),\{\Gamma_{K}(p)^{(v)}\}_{v\geqslant 0})$ is fully faithful. If $p\geqslant 3$ the proof of the equicharacteristic case in [4] uses explicit description of the ramification subgroups $\Gamma_{K}(p)^{(v)}$ modulo third commutators $C_{3}(\Gamma_{K}(p))\supset C_{p}(\Gamma_{K}(p))$. Then the mixed chracteristic case is deduced via application of the field-of-norms functor. In the case $p=2$ our approach does not work because modulo $C_{p}(\Gamma_{K}(p))=C_{2}(\Gamma_{K}(2))$ we are just in the abelian situation where the ramification filtration carries no essential information. However, in [7] we studied the images of ramification filtration in the Magnus algebra associated with $\Gamma_{K}(2)$ modulo subgroup $C_{3}(\Gamma_{K}(2))$ of third commutators. This new techniques allowed us to prove the case $p=2$ in the both (equicharacteristic and mixed characteristic) cases. Our method even worked in the case of higher local fields, cf. [5, 6].
@@
Question. How to describe the subgroups $\pi(\Gamma_{K}^{(v)})\subset GL_{\mathbb{Z}_{p}}(V)$?
@@ 0.1.4. Non-abelian explicit reciprocity laws (ERL)
Since $K_{\pi}/K$ is abelian all $\mathrm{Gal}(K_{\pi}/K)^{(v)}$ come from $\Gamma_{K}^{ab(v)}$ which are well-known by local class field theory. Then the corresponding images of $\Gamma_{K}^{(v)}$ can be found via one-or-another ERL for the Hilbert symbol. For example, suppose $K=k((t_{0}))$ contains a primitive $p^{M}$-th root of unity and $[K_{\pi}:K]=p^{M}$ then $K_{\pi}=K(\root p^{M} \of{v_{\pi}})$, $v_{\pi}\in K^{*}$. So the images $\pi(\Gamma_{K}^{(v)})$ can be calculated via calculations with the $M$-th Hilbert symbols $(u,v_{\pi})$, where $u\equiv 1\,\mathrm{mod}\,t_{0}^{v}$. The study of different ERL’s for the Hilbert symbol (and its analogs) is one of most popular areas of Algebraic Number Theory. It has very rich history [26] started with Kummer’s work in 1859. All known ERL’s have more-or-less similar structure: the answer appears as a residue of a suitable explicitly constructed differential form. As example, we can refer to fundamental contributions of Shafarevich [31], Artin-Hasse [14], Iwasawa [21], Brückner-Vostokov [16, 33], Witt [18], Coates-Wiles [17], et al. If $V$ from above Subsection 0.1.3 is not cyclic then one-or-another non-abelian version of the Hilbert symbol and the corresponding ERL’s could be used to study a similar question in full generality. Such symbol should use as essential ingredients nilpotent versions of local class field theory and appropriate versions of Kummer or Artin-Schreier theory. The idea of creating nilpotent class field theory was very popular for a very long time, e.g. [31, 23, 22]. Recently, F. Laubie [24] obtained such theory by generalizing the approach from [23]. It is based on the idea of iteration of Lubin-Tate theories via the field-of-norms functor. This theory posseses many nice properties but is not adjusted for obtaining (even very well-known) explicit information about involved Galois groups. A nilpotent analog of the Artin-Schreier theory was constructed by the author in [1, 2]; it allowed us to obtain new results about ramification filtration in [1, 3]. Introduction of nilpotent analogue of Hilbert symbol could be stated now as the question about linking the Laubie and our theories. This will allow us to benefit from functorial properties of the first and explicit constructions of the second theories. Notice that the case of “small” representations [13] gives a positive feeling about existence of ERL’s in the non-abelian context.
@@ 0.1.5. Geometrical approach to ramification filtration
The above applications not only illustrate a fundamental role of explicit results from [1, 2, 3] but also rise the question about their geometric origin and more structural approach. The proofs in [1] are reasonably technical but their modulo $p^{M}$-version in [2, 3] relies on heavy computations with enveloping algebras. When working with Lie algebras of nilpotent class $<p$ the most efficient and clear approach should use one-or-another deformation theory. Such theory was recently constructed in the case of Galois groups of period $p$ in [12] and this gave us a geometrical approach to [1]. In this paper we construct more general deformation theory which adjusted to the modulo $p^{M}$ situation. This theory not just gives the results of [2, 3] but also considerably simplifies the approach from [12]. We expect that this new techniques could be applied to develop the modulo $p^{M}$-case of results obtained in the mixed characteristic case in [9, 10]. Note that in [12] we gave also an explicit description of arbitrarily long segments of ramification filtration (not just of a single ramification subgroup) in the case of Galois groups of period $p$. This description admits a straight generalisation to the modulo $p^{M}$-situation. As a result, we have a substantial control on explicit information about arithmetical structure of Galois groups of cdvf’s. We also expect the existence of other fundamental relations between Arithmetic and Geometry of cdvf’s.
0.2. The content of the paper
Everywhere in the paper $M\in\mathbb{N}$ and $v_{0}\in\mathbb{R}_{>0}$ are fixed. For $s\in\mathbb{N}$ and a topological group $G$, we denote by $C_{s}(G)$ the closure of the subgroup of $G$ generated by commutators of order $\geqslant s$. If $L$ is a toplogical Lie algebra then $C_{s}(L)$ is the closure of the ideal generated by commutators of order $\geqslant s$. For any topological $A$-modules $M$ and $B$, we use the notation $M_{B}:=M\hat{\otimes}_{A}B$.
@@ 0.2. The content of the paper
Consider a decreasing filtration by ramification subgroups in upper numbering $\{\mathcal{G}^{(v)}_{<p,M}\}_{v\geqslant 0}$ of $\mathcal{G}_{<p,M}$. In [12] we constructed a deformation theory which allowed us to retrieve (in essentially geometrical way) the results from [1]. In this paper, we obtain a generalisation (and also a considerable simplification) of our approach from [12] to the case of Galois groups $\mathcal{G}_{<p,M}$ with arbitrary exponent $p^{M}$. We also explain how this theory can be applied to recover the structure of the ramification subgroups $\mathcal{G}_{<p,M}^{(v)}$. Our approach could be overviewed as follows.
a) Let $\mathcal{K}=k((t_{0}))$, where $k\simeq\mathbb{F}_{p^{N_{0}}}$. Then we can use the uniformiser $t_{0}$ to define the lifts modulo $p^{M}$, $O_{M}(\mathcal{K})=W_{M}(k)((t))\subset W_{M}(\mathcal{K})$ and $O_{M}(\mathcal{K}_{sep})=W_{M}(\sigma^{M-1}\mathcal{K}_{sep})[t]\subset W_{M}(\mathcal{K}_{sep})$. Here $W_{M}$ is the functor of Witt vectors of length $M$ and $t=[t_{0}]\in W_{M}(\mathcal{K})$ is the Teichmuller representative of $t_{0}$. This allows us to apply the nilpotent version of Artin-Shreier theory, cf. [2], to fix an identification $\eta^{0}:\mathcal{G}_{<p,M}\simeq G(\mathcal{L})$, where $\mathcal{L}$ is explicitly constructed pro-finite Lie $\mathbb{Z}/p^{M}$-algebra. The identification $\eta^{0}$ appears via choosing a suitable $e\in\mathcal{L}_{O_{M}(\mathcal{K})}$, taking $f\in\mathcal{L}_{O_{M}(\mathcal{K}_{sep})}$ such that $\sigma(f)=e\circ f$ and setting for any $\tau\in\mathcal{G}_{<p,M}$, $\eta^{0}(\tau)=(-f)\circ\tau(f)\in G(\mathcal{L})$. This provides us also with “ramification” ideals $\mathcal{L}^{(v)}$ in $\mathcal{L}$ such that $\eta^{0}(\mathcal{G}_{<p,M}^{(v)})=\mathcal{L}^{(v)}$.
b) The extension of scalars $\mathcal{L}_{W(k)}$ appears with the set of explicit generators $\{D_{an}\ |\ \mathrm{gcd}(a,p)=1,n\in\mathbb{Z}/N_{0}\}\cup\{D_{0}\}$. We use them to define a decreasing filtration of ideals $\mathcal{L}=\mathcal{L}(1)\supset\mathcal{L}(2)\supset\ldots\supset\mathcal{L}(p)$ and provide ¯$\bar{\mathcal{L}}=\mathcal{L}/\mathcal{L}(p)$ with induced filtration ¯$\bar{\mathcal{L}}(w)$, $1\leqslant w\leqslant p$. The algebra $\bar{\mathcal{L}}$ is considerably smaller than $\mathcal{L}$; it is generated over $W(k)$ by all $D_{an}$ with $0\leqslant a<(p-1)v_{0}$. However, we prove that $\mathcal{L}(p)\subset\mathcal{L}^{(v_{0})}$ and, as a result, it will be sufficient to describe the image ¯$\bar{\mathcal{L}}^{(v_{0})}$ of $\mathcal{L}^{(v_{0})}$ in $\bar{\mathcal{L}}$.
@@
Acknowledgements. The author is very grateful to the referee for encouraging comments, pointing out a few inexactitudes and suggesting several simplifications.